calculate the critical angle of a material having a refractive index of 1.52 when light pass from this material to air
1
Expert's answer
2019-08-26T10:42:39-0400

The critical angle


"\theta=\sin^{-1}\left(\frac{n_r}{n_i}\right)=\sin^{-1}\left(\frac{1.00}{1.52}\right)=41.1^{\circ}"